Circular No. 8853

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2007 J7 (LINEAR-MUELLER) Comet P/1998 S1 (cf. IAUC 7031) has been recovered independently

by R. H. McNaught on CCD images taken with the 0.5-m Uppsala Schmidt at Siding Spring on May 13 and by J. V. Scotti on CCD images taken with the Spacewatch 1.8-m f/2.7 reflector at Kitt Peak on June 26.5 UT. McNaught described the comet as faint and almost stellar, but with a 20′′ tail in p.a. 240^{o}. Scotti noted the comet to have a 9′′ coma and a 64′′ tail in p.a. 241^{o}. The indicated ΔT correction to the prediction on MPC 51824 is +0.03 day (ΔT = +0.013 day to the prediction by S. Nakano in the 2007 Comet Handbook, p. H9). The orbital elements by B. G. Marsden below are from 265 observations, 1998–2007 (mean residual 0′′.6).

2007 UT α_{2000} δ_{2000} Mag. Observer
May 13.78886 23^{h}22^{m}36.^{s}76 −9^{o}14′02′′.9 18.2 McNaught
June 26.45086 0 08 30.27 −3 24 01.9 18.1 Scotti
26.45843 0 08 30.69 −3 23 57.4 18.0  _{”}
26.46666 0 08 31.06 −3 23 53.7 18.0  _{”}

Epoch = 1998 Nov. 3.0 TT

T = 1998 Nov. 2.7434 TT  ω = 26.^{o}3621
e = 0.416704  Ω = 359.1837 }2000.0
q = 2.548366 AU  i = 10.5598

a = 4.368908 AU n^{o} = 0.1079306 P = 9.132 years Epoch = 2007 Dec. 6.0 TT

T = 2007 Dec. 16.1419 TT  ω = 26.^{o}4358
e = 0.415879  Ω = 359.1499 }2000.0
q = 2.552135 AU  i = 10.5462

a = 4.369188 AU n^{o} = 0.1079202 P = 9.133 years

COMETS C/2005 B5 AND C/2005 B6 (SOHO) Two additional Kreutz sungrazing comets (cf. IAUC 8850) have been

found by B. Zhou on 2005 archival SOHO website images. Each was stellar in appearance and faint, reaching mag ∼ 7–7.5.

Comet 2005 UT α_{2000} δ_{2000} Inst. F MPEC
C/2005 B5 Jan. 22.221 20^{h}33.^{m}1 −21^{o}40′ C3 BZ 2007-L16
C/2005 B6  30.571 21 06.8 −18 58 C3 BZ 2007-L16
2007 June 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green